OpenStudy (anonymous):

Expand (x+3)^6 using pascal's triangle

OpenStudy (slaaibak):

Sixth row of the triangle: 1 6 15 20 15 6 1 so it's x^6 + 6 * x^5 * 3 + 15 * x^4 * 3^2 + 20 * x^3 * 3^3 + 15 * x^2 * 3^4 + 6 * x^1 * 3^ 5 + 3^6

OpenStudy (slaaibak):

Simplified: x^6+18 x^5+135 x^4+540 x^3+1215 x^2+1458 x+729

myininaya (myininaya):

1 n=1 1 1 n=2 1 2 1 n=3 1 3 3 1 n=4 1 4 6 4 1 n=5 1 5 10 10 5 1 n=6 1 6 15 20 15 6 1 \[(x+y)^6=1 \cdot x^6y^0+6\cdot x^5y^1+15 \cdot x^4y^2+20\cdot x^3y^3+15\cdot x^2y^4+6 \cdot x^1y^5+1 \cdot x^0y^6\]

myininaya (myininaya):

replace y with 3

myininaya (myininaya):

that is y^6 at the end